Consider the following data set:

11 11 19 3 6 5 2 8 7 7 5 2

The value of the 3rd quartile is:

9

9.5

8

none of the above
Answer #1

Solution :

Arranging Observations in the ascending order,
2,2,3,5,5,6,7,7,8,11,11,19

Here, n=12

Q3 = (3(n+1)4)th value of the observation

= (3⋅134)th value of the observation

= (9.75)th value of the observation

= 9th observation +0.75[10th-9th]

= 8 + 0.75[11 - 8]

= 8 + 0.75( 3 )

= 8 + 2.25

= 10.25

3rd quartile = 10.25

None of the above